In a triangle, the measures of the second angle is 6 times the measure if the first angle. The measure of the third angle is 18 less than the sum of the measures of the second angle and the square of the first angle. Find the measure of each angle.
I tried this problem, but I got stuck after the 2nd equation which included finding the square of the first angle.

Answer by scott8148(6628)

f + s + t = 180 ___ t = 180 - f - s

s = 6f

t = f^2 + s - 18

180 - f - 6f = f^2 + 6f - 18 ___ 0 = f^2 + 13f - 18

use quadratic formula to find f. then substitute